NvidiaIntely (nvidiaintely):

Find the volume of a sphere with surface area equal to 100π ft2.

NvidiaIntely (nvidiaintely):

@OtherWorldly

NvidiaIntely (nvidiaintely):

@NerdyChick16, can you help me?

OpenStudy (otherworldly):

srry i had to go for a bit does the question tell u what the radius is

NvidiaIntely (nvidiaintely):

Nope :(

OpenStudy (nerdychick16):

would it be 4.19×106 ? if not I dunno :\

NvidiaIntely (nvidiaintely):

Hmmm....this is hard. Let me re-open this.

OpenStudy (otherworldly):

V= 4/3 pi r^3 A= 4 pi r^2 since we already know what the surface area is we just plug it in so 100 pi= 4 pi r^2 ignore the pi and divide by 4 and it will be 25= r^2 so the radius is 5 because 5 x 5 = 25 now that we know what the radius is plug it in the volume formula for sphere

OpenStudy (otherworldly):

so V= 4/3 pi 5^3 simplified: V= 4/3 pi 125 ignore the pi again and multiply 4/3 x 125

OpenStudy (otherworldly):

what would u get @NvidiaIntely